OpenStudy (anonymous):

Identify the x- and y-intercepts of f(x). Without graphing g(x), identify its x- and y-intercepts. f(x)=-3/2x+9;g(x)=2/3f(x)

OpenStudy (ash2326):

@michae1 You know what x and y intercepts mean?

OpenStudy (anonymous):

@ash2326 they are the x axis and y axis on a graph, correct?

OpenStudy (ash2326):



OpenStudy (ash2326):

The line where it cuts x axis, is the x intercept where it cuts y axis , is y intercept

OpenStudy (anonymous):

oh yea. i know this. im just doing this problem on the web cause i don't really understand this way of graphing

OpenStudy (ash2326):

f(x)=-3/2x+9 to find x intercept put f(x) ( or y) =0 and find value of x to find y intercept put x=0 and find value of f(x) Can you try?

OpenStudy (anonymous):

these are the possible answers... a) f(x): 6, 9 g(x): 6, 6 b) f(x): 6, 9 g(x): 6, 9 c) f(x): 6, 9 g(x): 4, 6 d) f(x): 6, 9 g(x): 4, 9

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

x=6?

OpenStudy (ash2326):

yes, good :) now find y intercept

OpenStudy (anonymous):

i dont get how to solve g(x)=2/3f(x)

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

okay i need help..

OpenStudy (ash2326):

We need to find g(x) \[g(x)=\frac 2 3 f(x)\] \[g(x)=\frac 2 3 \times(\frac {-3}{2} x +9)\] multiply and find g(x)

OpenStudy (anonymous):

i got 9

OpenStudy (ash2326):

for y intercept of f(x)?

OpenStudy (anonymous):

no for g(x)...

OpenStudy (ash2326):

Try again

OpenStudy (anonymous):

i divided negative one to negative 9

OpenStudy (anonymous):

2/3×(−3/2x

OpenStudy (anonymous):

what dos that =

OpenStudy (ash2326):

\[g(x)=\frac 2 3 \times(\frac {-3}{2} x +9)\] \[=\frac 2 3 \frac 3 2(-x)+\frac 2 3 \times 9\] \[=-x+6\]

OpenStudy (anonymous):

so then we plug in 6?

OpenStudy (ash2326):

for x intercept put g(x)=0 y intercept x=0

OpenStudy (anonymous):

so g(x):6, 6

OpenStudy (ash2326):

yes :) do you understand this?

OpenStudy (anonymous):

yea i do now

OpenStudy (ash2326):

cool :D